The ball is initially at 2ȋ + 0ĵ + 3k̂ then it moved to 4ȋ + 0ĵ + 5k̂ , its total displacement is 6ȋ + 0ĵ + 8k̂ .


1
Expert's answer
2021-10-03T13:23:05-0400

If the body is initially at 2ȋ + 0ĵ + 3k̂ and then it moved to 4ȋ + 0ĵ + 5k̂, the total displacement is


"d=\sqrt{(r_{x2}-r_{x1})^2+(r_{y2}-r_{y1})^2+(r_{z2}-r_{z1})^2}=\\
=\sqrt{(4-2)^2+(0-0)^2+(5-3)^2}=2.83."